# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

### v3

@@ 3. CSBASG Construction
Intuitively, to create an ASG following Definition 3, we first need to do a pre-order traversal on the AST to identify and combine the semantically identical nodes. We start by constructing an isomorphism between $\mathcal{V}$ and $N\cup\Sigma$, then rewrite each AST link into an ASG edge. The isomorphism is trivial since AST nodes with the same word are combined into the same node in the ASG. Nonetheless, constructing the ASG edge set by the AST relations $\xi$ requires a cautious step: merging two AST nodes requires clarifying the orders of the visit to avoid confusion. or instance, in the example given in Figure 1, if we consider the binary operator ‘‘IN’’ as a single node without specifying the order of execution, we would not specify which could be put at the right of the operator, ‘‘p.Tutors’’, ‘‘p.Tutors.Tutors’’ or ‘‘p.Tutors.Tutors.Tutors’’. A naive approach that deviates a bit from the ASG definition is to use two numbers for the weight values $w_{ij}$ instead of one, that is, for any of the existing link $x_{2}\in\xi(x_{1})$ in the AST, defining $v_{1}=\sigma(x_{1}),v_{2}=\sigma(x_{2})$. Then, we create an edge between $v_{1},v_{2}$ with a weight-pair $(\omega,t)\in\mathbb{N}\times\mathbb{N}$, specifying that $x_{2}$ is the $\omega$-th child of $x_{1}$ in the order defined for $\xi(x_{1})$, and it is the $t$-th time to visit $v_{1}$ in the program logic.
Nevertheless, we expect that with a matrix form of ASG, we could take advantage of their linearity, especially for the potential application of machine learning operations. In addition, a matrix form of the graphs could be more easily compared with respect to multiple implementations of a predicate or other code segments of interest. With the naive construction above, it is challenging to proceed since the mapping from the pair given above to the edge value that serves as an entry in the adjacency matrix is nontrivial. Here, we could define $\mathcal{M}:\mathbb{N}\times\mathbb{N}\to\mathbb{R}$ as mapping such that $\mathcal{M}(\omega,k)$ gives the entry in the matrix form of an ASG if it is the sole AST link in the ASG. The mapping could be defined arbitrarily, provided that it maintains injectivity to avoid the simple, different-edge-same-value confusion (we will call it Type 1 Confusion later). Such forms the real part (magnitude) of the edges in the CSBASG.
@@
Algorithm 1 is a high-level process to convert an AST to a CSBASG. It follows the process of an abstract interpretation, where $\vec{t}$ in the recursive part gives a simplified interpretation state tracking the times of node visitations in the pre-order traversal of the AST. We initialize the nodemap and the list of angular signatures for each node, with the only element being the dummy overall root, and assign each node an angular signature with regard to the predefined assignment, outlining every unique signature; the number of unique angular signatures (semantically distinct nodes) is the number of rows and columns of the resulting adjacency matrix. After that, we initialize a state vector $\vec{t}$ to count the times of the order of the visit to each node. A recursive function will then iterate over the AST beginning from the overall root, update the adjacency matrix for each downlink from the root node according to the composition function $\beta$ and the encoding method $\mathcal{M}$, while updating the counter of the given node with the corresponding entry of the state vector signaling the pre-order time of identical appearance in the AST before recursively visit the subtree with the new root being a child of the local root.
In general, with a static signature assignment $\varphi$, the algorithm terminates in $O(|X|)$ since each of the nodes in the AST was visited exactly once and $|\mathcal{V}|<|X|$ strictly. We say an ASG is generated by an AST if the AST was converted to the same ASG with the same signature assignment and magnitude mapping.
@@
By Definition 5, for a given Alloy predicate as an input, define a fixed unique root $r\equiv v_{1}$, which is the predicate itself, and such root node is unique in each predicate. We can intuitively assign $\theta_{1}=\varphi(\sigma(r))\equiv 0$ as a starting point. Besides, we expect the same input parameters for the pairs of predicates to be directly compared, so the overall root node of each predicate only has one child, the local root expression or formula. Therefore, the only work left is to define the signature assignment function $\varphi$ on the subset that could present under a tree rooted by an expression or formula.
Another notable function, $\mathcal{M}$, could be defined in multiple ways if it maintains a one-on-one correspondence between the syntactic positions and the execution or tree-walking order. However, for an AST with multiple syntactically and semantically identical nodes, multiple edges connect them, but they could have different positions $\omega$ and thus different $\mathcal{M}$ values. So, given an AST, we can define the three types of confusion that could happen in an ASG construction:
@@
Here, we introduce the polynomial-based, integer-magnitude static CSBASG encoding for a theoretical language: each AST generated from its CFG has a maximum number of children $p$. Then, let $T$ be the maximum number of syntactic and semantically identical nodes in the code segment. For the AST link drawn from a node to its $(\omega+1)$-th child, let $\mathcal{M}(\omega,t)=2^{(t-1)p+\omega}$ and $\beta$ be a simple summation of the $\mathcal{M}$ values. Our construction of the scheme was defined in the order of $O(2^{pT})$. It could be despairing at a glance, yet consider that in many languages without a list expression, $p\leq 5$ could be achieved, and $T$ only increases with the density and size of the code chunk, which raw AST representation also suffers such a polynomial growth.
Lemma 0 ().
@@
By Lemma 2, the summation-of-exponentials gives the most compact and accurate solution for the complete enumeration of the graph, and the creation of such a graph also depends on the enumerability of the grammar. However, a representation within the 32- or 64-bit bound could be achieved for less scaled language, such as Alloy. A heuristic generated by representation learning could also be achieved by learning the logarithmic parameters for a more generally purposed language with more syntactic elements and a larger AST.
@@
The decoding process is straightforward in theory. Words of the code are stored in correspondence to rows and columns of the matrix; each row or column corresponds to an individual word, and vice versa. Formally, we designate $L:\mathcal{G}\to\mathcal{T}$ as the parser that translates an ASG back to its uniquely corresponding AST. $L$ is therefore a counterpart of Algorithm 1 and bounded by the constructions of $\mathcal{M}$ and $\beta$ functions. For each of the nonzero entries in the ASG adjacency matrix, we can break it into concrete links by reversing $\mathcal{M}$ and $\beta$. For the binary polynomial-based magnitude encoding given above, we take the integer value of the magnitude $|a_{ij}|$ and break it with regard to its binary form: each digit signing $1$ in the binary form corresponds to a concrete AST link. In this case, an ASG node was broken into a set of AST nodes, each with its corresponding outer edges by the visit order. Since the possible edges are all examined once and the size of the entries determines the binary forms, we have a complexity of $O(|\mathcal{E}|\log(\max(w_{i,j})))$ or $O(|\mathcal{V}|^{2}\log(\max(w_{i,j})))$ for the decoding process.
@@ 4. Application of CSBASG on Alloy
Alloy, a specification language, has plenty of features that construct a friendly scenario for CSBASG to apply. Firstly, as a first-order logic-based declarative language, we could easily construct a finite set of symbols given a model source code set. Besides, there are few infinitely extensible structures in its grammar, which limits the usage of empirical estimates of the maximum nodes under a given structure; furthermore, even the infinitely extensible structures could be assumed with a relatively low count of children. Its simplicity makes an ideal starting point for CSBASG representation.
@@
To begin with, we utilize a parser that classifies the syntactic properties of the code keywords into a limited number of categories. After that, we need to set a $p$ value corresponding to the maximum number of children for each category, for instance, $1$ for unary operators, $2$ for binary operators, $3$ for if-then-else, and a more significant number for the several uncovered infinite categories. This could cause ambiguity in some edge cases. Still, since Alloy was used in modeling the structures within the scope of a category of sets, there will unlikely be any long enumerations over 10 or 15 items, so we use $p=17$ for the categories without a precise, finite maximum children number. While completeness is not guaranteed, we are aiming at solving the absolute most of the cases; while it is possible to break the enumerations by putting the list of items into a linked structure, in most cases, the tradeoff of completeness in the edge cases could deliver more benefits in intuitiveness and compactness of the representation.
@@
Inside each category, the nodes are differentiated by their semantic contents, which are still finitely enumerable since we have access to the set signatures defined within a model file, and there are almost no numerical constants, so it is safe to assume that constants are limited to booleans and small integers. In our approaches, we divided the AST nodes under a paragraph of an expression or formula into 19 categories, plus the zero-signature node signing the local root. Devolved from the syntactic categories are semantic subcategories, and the number of subcategories within a syntactic category ranges from 1 for the simple, invariant nodes to 32, which is the number of possible binary expressions.
@@
Finally, we are required to assign the angular signatures for each node. In our practice, we used a simple encoding by dividing the interval of a circle $(-\pi,\pi)$ into equal-length intervals with $D=\pi/13$ as the length of each interval. For each interval, we assign subticks that divide it into 128 or 65536 sectors, depending on the syntactic category. By the scheme above, we have obtained each AST node’s syntactic-semantic pair $(\Psi,\psi)$ using the results generated from the parser. Then, we map them on the unit circle as their signatures: each syntactic category goes with the beginning of a $\pi/13$-length interval, and each subtick above the interval starting point corresponds to a semantic subcategory. Note that this is just a demonstration of encoding and is only defined for convenience but not learned for optimality, and the angular correspondence could be any bijective mappings.
@@ 5. Experiment
Figure 4. An example of a pair of student-written mutant predicate (Inv2) and the ground truth (Inv2C). The ASG edges in form $(x,y)$ indicate the $x$-th visit of the sourcing node, and the positional relationship between the nodes is $y$. Black edges are present in both predicates, red edges are present in the mutant only, and blue edges are present in the ground truth only; a fixing process, in theory, removes all red edges and replaces them with blue ones.
Dataset #Models Mut.Type MutantCorrect%
lts 61 CORRECT 36.66 39.19
lts 113 UNDER 24.65 40.52
@@
classroom_fol 495 CORRECT 32.83 14.67
classroom_fol 223 UNDER 30.48 16.86
@@
cv_v1 52 CORRECT 31.73 39.83
cv_v1 118 UNDER 36.05 41.48
@@
train 75 CORRECT 29.44 19.43
train 170 UNDER 23.47 25.23
train 277 BOTH 28.18 28.54
train 174 UNDER 32.70 24.43
trash_rl 252 CORRECT 20.25 6.73
trash_rl 150 UNDER 13.42 7.10
trash_rl 267 BOTH 14.85 6.99
trash_rl 62 UNDER 31.67 11.68
cv_v2 31 CORRECT 30.95 40.47
cv_v2 40 UNDER 38.28 43.79
cv_v2 21 BOTH 38.37 43.08
cv_v2 12 UNDER 29.16 36.21
classroom_rl 495 CORRECT 32.83 14.67
classroom_rl 223 UNDER 30.48 16.86
@@
production 25 CORRECT 14.75 21.12
production 25 UNDER 18.84 20.73
@@
Table 1. Compactness test results of the dataset. Each of the sub-dataset represents a category of problems within Alloy4Fun(alloy4fun, 21), and the student-written mutant predicates are categorized into four mutant types given their coverage of the set indicated by the ground truth: CORRECT, OVERconstrained, UNDERconstrained, or BOTH over- and underconstrained. The percentages given are the ratio of the reduction of nodes in the ASG representation compared to the raw AST.
5.1. Dataset
Our dataset encompasses 6,307 models from Alloy4Fun(alloy4fun, 21), sourced from 8 different categories of problem, filtered from a larger dataset to ensure each is compliable, runnable, nonempty, and the student-written model is not completely identical with the ground truth. Each model has two predicates under a set of definitions of relations; a student writes one named InvX (or PropX, X is an identifying number) and is potentially faulty as a mutant predicate, and another called InvXC is the ground truth that gives a correct formula. Then we run two predicates, overconstrained and underconstrained, which determine if there are cases that satisfy InvX but not InvXC, or vice versa. If a model has no cases for both overconstrained and underconstrained predicates, then InvX is formally correct. This gives four categories of the models: correct ($InvX=InvXC$), overconstrained only ($InvX\subsetneq InvXC$), underconstrained ($InvXC\subsetneq InvX$), and both overconstrained and underconstrained ($InvX\not\subset InvXC\land InvXC\not\subset InvX$), covers all possible cases.
@@ 5.2. Metric 1: Compactness of representation
We evaluate the compactness of the representation by the percentage of nodes in the ASG representation that are less than the raw AST. Overall, we achieved an impressive 27.25% reduction of nodes over the dataset, considering there is no loss of information, and the performance for each category of predicates is given in Table 1.
@@
Note that since the dataset is mostly single predicates designed as after-class practices for college students, the probability of semantic identical nodes in the AST is relatively low, and for a sub-AST with a large scale, the performance would increase accordingly in the single-metric of node reduction. Moreover, the size of the adjacency matrix scales up with $O(N^{2})$ with the number of nodes, so a denser matrix could be achieved with the ASG scheme.
@@
However, this metric should not be confused with saving spaces. Since in an AST the entries in the adjacency matrix are either 0 (indicating no links) or 1 (indicating a direct link), which only occupy 1-bit per entry in an optimized language or computational method, but Lemma 2 indicates that there is no complete encoding without information loss that takes a number smaller than $O(2^{pT})$ which occupies $pT$-bits at least, so space saving is neither guaranteed or indicated by the inference and experiment above.
5.3. Metric 2: A case study of comparability
Comparing a pair of code snippets to train a model for auto-correction could be tedious, especially for the syntactically correct faulty predicates. Comparability of the CSBASG could point out a direction of inquiry by the graph mutations. We have a simple formula to decode its multiedges, given our construction of the CSBASG with the polynomial method.
Since the dataset was written by students who have little experience in Alloy modeling and some of them write models that have few, if not zero, common edges; however, among them, there could also be some manual repairs that could be shown as a process of applying a set of ‘‘Atomic Mutations’’. We call an atomic mutation either adding or removing an edge in the ASG; a relationship change between two nodes could also be represented as removing the old edge and adding the new one.
Figure 4 gives an example of the decoded CSBASG construction for comparing two predicates in the same declarative environment, the faulty and correct solution to the same problem. The common parts, as reflected by the subgraph characterized with black edges, show the correct declaration of the first two variables by the student given the right part of the graph and the trivial correctness of the fields associated with the variables as elements of the pre-declared sets were shown in the bottom-left corner. However, most of the graph shows significant deviation of logic, as expected for the student-written mutant being both overconstrained and underconstrained. Straightforwardly, each of the red or blue edges in the graph is an atomic mutation between the pair of predicates; the red edge is a link to be broken, and the blue edge is a link to be established in a hypothetical automatic fixing application. In the example above, we could see that out of 41 edges shown in Figure 4, there are 10 common edges presenting in both the faulty model and the ground truth, 20 edges are added and 11 edges are removed from the faulty model, which are performs of atomic mutations.
[…]
Dataset #Models Mut.Type MutantCorrect%
lts 61 CORRECT 51.61 49.31
lts 113 UNDER 69.66 31.03
@@ 5.3. Metric 2: A case study of comparability
classroom_fol 495 CORRECT 82.07 71.77
classroom_fol 223 UNDER 77.54 63.51
@@
cv_v1 52 CORRECT 62.88 38.86
cv_v1 118 UNDER 58.99 34.84
@@
train 75 CORRECT 66.66 63.86
train 170 UNDER 70.06 49.00
train 277 BOTH 75.16 50.97
train 174 UNDER 73.29 60.55
trash_rl 252 CORRECT 90.64 73.49
trash_rl 150 UNDER 84.08 61.93
trash_rl 267 BOTH 89.39 62.35
trash_rl 62 UNDER 90.21 75.24
cv_v2 31 CORRECT 63.93 38.38
cv_v2 40 UNDER 65.78 33.73
cv_v2 21 BOTH 67.67 35.04
cv_v2 12 UNDER 70.70 36.82
classroom_rl 495 CORRECT 82.07 71.77
classroom_rl 223 UNDER 77.54 63.51
@@
production 25 CORRECT 68.12 49.40
production 25 UNDER 71.80 45.60
@@
Table 2. Comparability quantification results in the per-category experiment. The percentages given are the ratio of the edges that also presenting in its mutant or corrected counterpart with the same relationship (position and execution order), that is the percentage of the shared edges in each predicate.
6. Future Works

### v4

@@ 3. CSBASG Construction
Intuitively, to create an ASG following Definition 3, we first need to do a pre-order traversal on the AST to identify and combine the semantically identical nodes. We construct an isomorphism between $\mathcal{V}$ and $N\cup\Sigma$, then rewrite each AST link into an ASG edge. The isomorphism is trivial since AST nodes with the same word are combined into the same node in the ASG. Nonetheless, constructing the ASG edge set by the AST relations $\xi$ requires a cautious step: merging two AST nodes requires clarifying the orders of the visit to avoid confusion. For instance, in the example given in Figure 1, if we consider the binary operator “IN” as a single node without specifying the order of execution, we would not specify which of the following would be put as the right-hand operand: “p.Tutors”, “p.Tutors.Tutors” or “p.Tutors.Tutors.Tutors”. A naive approach that deviates a bit from the ASG definition is to use two numbers for the weight values $w_{ij}$ instead of one, that is, for any of the existing link $x_{2}\in\xi(x_{1})$ in the AST, defining $v_{1}=\sigma(x_{1}),v_{2}=\sigma(x_{2})$. Then, we create an edge between $v_{1},v_{2}$ with a weight-pair $(\omega,t)\in\mathbb{N}\times\mathbb{N}$, specifying that $x_{2}$ is the $\omega$-th child of $x_{1}$ in the order defined for $\xi(x_{1})$, and it is the $t$-th time to visit $v_{1}$ in the program logic.
Nevertheless, we expect that for a matrix form of the ASG, we could take advantage of their linearity, especially for the potential application to machine learning operations. In addition, a matrix form of the graphs could be more easily compared with respect to multiple implementations of a predicate or other code segments of interest. With the naive construction above, it is challenging to proceed since the mapping from the pair given above to the edge value that serves as an entry in the adjacency matrix is nontrivial. Here, we could define $\mathcal{M}:\mathbb{N}\times\mathbb{N}\to\mathbb{R}$ as mapping such that $\mathcal{M}(\omega,k)$ gives the entry in the matrix form of an ASG if it is the sole AST link in the ASG. The mapping could be defined arbitrarily, provided it maintains injectivity to avoid the simple, different-edge-same-value confusion (we will call it Type 1 Confusion later). Such forms the real part (magnitude) of the edges in the CSBASG.
@@
Algorithm 1 gives the high-level process to convert an AST to a CSBASG. It follows the process of an abstract interpretation, where $\vec{t}$ in the recursive part gives a simplified interpretation state that tracks the times of node visitations in the pre-order traversal of the AST. We first initialize the nodemap and the list of angular signatures for each node, with the only element being the dummy overall root. Then, we assign each node an angular signature with regard to a predefined assignment, which outlines every unique signature. The number of unique angular signatures (semantically distinct nodes) is the number of rows and columns of the resulting adjacency matrix. After that, we initialize a state vector $\vec{t}$ to count the times of the order of visitation to each node. A recursive function then iterates over the AST beginning from the overall root, updates the adjacency matrix for each downlink from the root node according to the composition function $\beta$ and the encoding method $\mathcal{M}$, while updating the counter of the given node with the corresponding entry of the state vector signaling the pre-order time of identical appearance in the AST before recursively visiting the subtree with the new root being a child of the local root.
In general, with a static signature assignment $\varphi$, the algorithm terminates in $O(|X|)$ since each of the nodes in the AST is visited exactly once and $|\mathcal{V}|<|X|$ strictly. We say an ASG is generated by an AST if the AST was converted to the same ASG with the same signature assignment and magnitude mapping.
@@
By Definition 5, for a given Alloy predicate as an input, we define a fixed unique root $r\equiv v_{1}$, which is the predicate itself, and such root node is unique in each predicate. We can intuitively assign $\theta_{1}=\varphi(\sigma(r))\equiv 0$ as a starting point. Besides, we expect the same input parameters for the pairs of predicates to be directly compared, so the overall root node of each predicate only has one child, the local root expression or formula. Therefore, the only work left is to define the signature assignment function $\varphi$ on the subset that could be present under a tree rooted by an expression or formula.
The function, $\mathcal{M}$, could be defined in multiple ways if it maintains a one-on-one correspondence between the syntactic positions and the execution or tree-walking order. However, for an AST with multiple syntactically and semantically identical nodes, there can be multiple edges that connect them, but those edges could have different positions $\omega$ and thus different $\mathcal{M}$ values. So, given an AST, we can define the three types of confusion that could happen in an ASG construction:
@@
Here, we introduce a polynomial-based, integer-magnitude static CSBASG encoding for a theoretical language: each AST generated from its CFG has a maximum number of children $p$. Then, let $T$ be the maximum number of syntactic and semantically identical nodes in the code segment. For the AST link drawn from a node to its $(\omega+1)$-th child, let $\mathcal{M}(\omega,t)=2^{(t-1)p+\omega}$ and $\beta$ be a simple summation of the $\mathcal{M}$ values. Our construction of the scheme was defined in the order of $O(2^{pT})$. It could be despairing at a glance, yet consider that in many languages without a list expression, $p\leq 5$ could be achieved, and $T$ only increases with the density and size of the code chunk, and raw AST representations also suffers such a polynomial growth.
Lemma 0.
@@
By Lemma 2, the summation-of-exponentials gives the most compact and accurate solution for the complete enumeration of the graph, and the creation of such a graph also depends on the enumerability of the grammar. However, a representation within the 32- or 64-bit bound could be achieved for a less-scaled language like Alloy. A heuristic generated by representation learning could also be achieved by learning the logarithmic parameters for a more general-purpose language with more syntactic elements and a larger AST.
@@
The decoding process is straightforward in theory. Words of the code are stored in correspondence to rows and columns of the matrix; each row or column corresponds to an individual word, and vice versa. Formally, we designate $L:\mathcal{G}\to\mathcal{T}$ as the parser that translates an ASG back to its uniquely corresponding AST. $L$ is therefore a counterpart of Algorithm 1 and bounded by the constructions of $\mathcal{M}$ and $\beta$ functions. For each of the nonzero entries in the ASG adjacency matrix, we can break it into concrete links by reversing $\mathcal{M}$ and $\beta$. For the binary polynomial-based magnitude encoding given above, we take the integer value of the magnitude $|a_{ij}|$ and break it with regard to its binary form: each digit signing $1$ in the binary form corresponds to a concrete AST link. In this case, an ASG node is broken into a set of AST nodes, each with its corresponding outer edges by the visit order. Since the possible edges are all examined once and the size of the entries determines the binary forms, we have a complexity of $O(|\mathcal{E}|\log(\max(w_{i,j})))$ or $O(|\mathcal{V}|^{2}\log(\max(w_{i,j})))$ for the decoding process.
@@ 4. Application of CSBASG on Alloy
There are two main features of Alloy that make for an ideal starting point to apply our CSBASG representation. First, as a first-order logic-based declarative language, we can easily construct a finite set of symbols given a model’s source code. Second, while there are a few infinitely extensible structures in its grammar, which does limit the usage of empirical estimates for the maximum nodes under a given structure, even the infinitely extensible structures could be assumed to have a relatively low count of children.
@@
To begin, we utilize a parser that classifies the syntactic properties of the code keywords into a limited number of categories. After that, we need to set a $p$ value corresponding to the maximum number of children for each category, for instance, $1$ for unary operators, $2$ for binary operators, $3$ for if-then-else, and a more significant number for the several uncovered infinite categories. This could cause ambiguity in some edge cases. Still, since Alloy is used to model structures within a bounded scope for each category of sets, there will unlikely be any long enumerations over 10 or 15 items, so we use $p=17$ for the categories without a precise, finite maximum children number. While completeness is not guaranteed, we aim to solve completeness for the majority of cases. While it is possible to break the enumerations by putting the list of items into a linked structure, in most cases, the tradeoff of completeness in the edge cases delivers more benefits in the intuitiveness and compactness of the representation.
@@
Inside each category, the nodes are differentiated by their semantic contents, which are still finitely enumerable since we have access to the signatures defined within a model file, and there are almost no numerical constants. Moreover, the numerical constant will be bound by the scope. Therefore, it is safe to assume that constants are limited to booleans and small integers. In our approaches, we divided the AST nodes under a paragraph of an expression or formula into 19 categories, plus the zero-signature node capturing the local root. Devolved from the syntactic categories are semantic subcategories, and the number of subcategories within a syntactic category ranges from 1 for the simple, invariant nodes to 32, which is the number of possible binary expressions.
@@
Finally, we need to assign the angular signatures for each node. In our practice, we used a simple encoding by dividing the interval of a circle $(-\pi,\pi)$ into equal-length intervals with $D=\pi/13$ as the length of each interval. For each interval, we assign subticks that divide it into 128 or 65536 sectors, depending on the syntactic category. By the scheme above, we have obtained each AST node’s syntactic-semantic pair $(\Psi,\psi)$ using the results generated from the parser. Then, we map them on the unit circle as their signatures: each syntactic category goes with the beginning of a $\pi/13$-length interval, and each subtick above the interval starting point corresponds to a semantic subcategory. Note that this is just a demonstration of encoding and is only defined for convenience but not learned for optimality, and the angular correspondence could be any bijective mappings.
@@ 5. Experiment
We address the following research questions:
  • RQ1: What degree of compactness does the CSBASG representation of an Alloy model achieve?
  • RQ2: Does our CSBASG representation of an Alloy model enable comparison between different predicates?
Table 1. Overview of problem set complexity.
Problem #Sig #Rel #Exe #AST
classroom_fol 5 3 15 10.00
classroom_rl 5 3 15 10.00
cv_v1 5 4 4 19.75
cv_v2 5 4 4 21.75
lts 3 1 7 19.71
production 5 3 4 14.25
train 6 3 18 23.44
trash_rl 3 1 10 4.80
5.1. Dataset
Our dataset encompasses 6,307 models from Alloy4Fun(Macedo et al., 2019), sourced from 8 different problem sets, filtered from a larger dataset to ensure each is compliable, runnable, nonempty, and the student-written model is not completely identical to the ground truth. For each problem set, Alloy4Fun has a base model that outlines all the signatures and relations, as well as a collection of empty predicates with English descriptions. Users can then attempt to fill in the predicate to match the English description, and their submission is checked against a backend oracle for correctness. Table 1 gives an overview of the complexity of each problem set: Column #Sig is the number of signatures, #Rel is the number of relations, #Exe is the number of different predicates uses can try to complete on Alloy4Fun, and #AST is the average number of AST nodes in the oracle solutions.
Our 6,307 models contain the base signatures from the Alloy4Fun problem set along with two predicates: a student-written predicate named InvX (or PropX, X is an identifying number) and another predicate called InvXC that is oracle solutions from Alloy4Fun. For each model, we run two predicates, overconstrained and underconstrained, which determine if there are cases that satisfy InvX but not InvXC, or vice versa. If a model has no cases for both overconstrained and underconstrained predicates, then InvX is formally correct. This gives four categories of the models: correct ($InvX=InvXC$), overconstrained only ($InvX\subsetneq InvXC$), underconstrained ($InvXC\subsetneq InvX$), and both overconstrained and underconstrained ($InvX\not\subset InvXC\land InvXC\not\subset InvX$), covers all possible cases.
Table 2. Compactness test results of the dataset.
Problem #Models Mut.Type Mut.Oracle%
lts 113 OVER 24.65 40.52
@@ 5.1. Dataset
lts 61 CORRECT 36.66 39.19
trash_rl 150 OVER 13.42 7.10
trash_rl 267 BOTH 14.85 6.99
trash_rl 62 UNDER 31.67 11.68
trash_rl 252 CORRECT 20.25 6.73
train 170 OVER 23.47 25.23
train 277 BOTH 28.18 28.54
train 174 UNDER 32.70 24.43
train 75 CORRECT 29.44 19.43
classroom_fol 223 OVER 30.48 16.86
@@
classroom_fol 495 CORRECT 32.83 14.67
cv_v2 40 OVER 38.28 43.79
cv_v2 21 BOTH 38.37 43.08
cv_v2 12 UNDER 29.16 36.21
cv_v2 31 CORRECT 30.95 40.47
cv_v1 118 OVER 36.05 41.48
@@
cv_v1 52 CORRECT 31.73 39.83
classroom_rl 223 OVER 30.48 16.86
@@
classroom_rl 495 CORRECT 32.83 14.67
production 25 OVER 18.84 20.73
@@
production 25 CORRECT 14.75 21.12
@@ 5.2. Metric 1: Compactness of Representation
We evaluate the compactness of the representation by comparing the number of nodes in the ASG representation compared to the number of nodes in the AST. Table 2 gives an overview of these results. Column Problem displays the problem set from Alloy4Fun (Macedo et al., 2019), and column #Models displays the number of unique student submissions. From each problem set, the student-written submissions, which we refer to as mutant predicates, are categorized further into four mutant types given their coverage of the set indicated by the ground truth: CORRECT, OVERconstrained, UNDERconstrained, or BOTH over- and underconstrained (Column Mut.Type). The percentages given are the ratio of the reduction of nodes in the ASG representation compared to the raw AST for the student submission (Column Mut.%) and the oracle solution (Column Oracle%), e.g., there are 32.83% fewer nodes in the ASG of correct classroom_fol student submissions than the AST.
@@
Overall, we achieved an impressive 27.25% reduction of nodes over the dataset with no loss of information. Interestingly, a smaller oracle does not guarantee less reduction. While the model with the smallest oracles by number of AST nodes, trash_rl, does in fact see the smallest reduction, the train model has the largest oracle but three models have larger reductions in nodes. All told the reduction the CSBASG provides over the AST is tied to the likelihood that the formula to be expressed has redundancy within its structure. Note that since the dataset is mostly single predicates designed as after-class practices for college students, the probability of semantic identical nodes in the AST is relatively low, and for a sub-AST with a large scale, the performance would increase accordingly in the single-metric of node reduction. Moreover, the size of the adjacency matrix scales up with $O(N^{2})$ with the number of nodes so that a denser matrix could be achieved with the ASG scheme.
Figure 5. An example of a pair of student-written mutant predicate (Inv2) and the ground truth (Inv2C). The ASG edges in form $(x,y)$ indicate the $x$-th visit of the sourcing node, and the positional relationship between the nodes is $y$. Black edges are present in both predicates, red edges are present in the mutant only, and blue edges are present in the ground truth only.
@@
However, this metric should not be confused with saving spaces. Since in an AST the entries in the adjacency matrix are either 0 (indicating no links) or 1 (indicating a direct link), which only occupy 1-bit per entry in an optimized language or computational method, but Lemma 2 indicates that there is no complete encoding without information loss that takes a number smaller than $O(2^{pT})$ which occupies $pT$-bits at least, so space saving is neither guaranteed nor indicated by the inferences and experiment above.
Table 3. Comparability quantification results of the dataset.
Problem #Models Mut.Type Mut.Oracle%
lts 113 OVER 69.66 31.03
@@
lts 61 CORRECT 51.61 49.31
trash_rl 150 OVER 84.08 61.93
trash_rl 267 BOTH 89.39 62.35
trash_rl 62 UNDER 90.21 75.24
trash_rl 252 CORRECT 90.64 73.49
train 170 OVER 70.06 49.00
train 277 BOTH 75.16 50.97
train 174 UNDER 73.29 60.55
train 75 CORRECT 66.66 63.86
classroom_fol 223 OVER 77.54 63.51
@@
classroom_fol 495 CORRECT 82.07 71.77
cv_v2 40 OVER 65.78 33.73
cv_v2 21 BOTH 67.67 35.04
cv_v2 12 UNDER 70.70 36.82
cv_v2 31 CORRECT 63.93 38.38
cv_v1 118 OVER 58.99 34.84
@@
cv_v1 52 CORRECT 62.88 38.86
classroom_rl 223 OVER 77.54 63.51
@@
classroom_rl 495 CORRECT 82.07 71.77
production 25 OVER 71.80 45.60
@@
production 25 CORRECT 68.12 49.4
5.3. Metric 2: Comparability of Representation
Comparing a pair of code snippets to train a model for automated correction could be tedious, especially for syntactically correct but faulty predicates. Therefore, the CSBASG representation’s comparability could enable graph mutations to annotate the pairs and create training data. We will call an atomic mutation either adding or removing an edge in the ASG. A relationship change between two nodes could also be represented as removing the old edge and adding the new one. To explore this idea, we have implemented a simple formula to decode the multiedges created when constructing the CSBASG using the polynomial method.
[…]
To take a closer look at comparability, Figure 5 gives an example of the decoded CSBASG construction for comparing a pair of predicates in the same declarative environment, which, in our case, is the faulty and correct solution to the same problem. The common parts, depicted as black edges, show the correct declaration of the first two variables by the student given the right part of the graph, and the trivial correctness of the fields associated with the variables as elements of the pre-declared sets are shown in the bottom-left corner. However, most of the graph shows significant deviation of logic, as expected since the student predicate is both overconstrained and underconstrained. Straightforwardly, each of the red or blue edges in the graph is an atomic mutation between the pair of predicates; the red edge is a link to be broken, and the blue edge is a link to be established in a hypothetical automatic fixing application. In the example above, we could see that out of 41 edges shown in Figure 5, there are 10 common edges, 20 edges are added, and 11 edges are removed from the faulty model, which outlines the atomic mutations.
6. Future Work
In this section, we present several future improvements to the CSBASG representation and planned applications.